when you reverse the digits in a certain two digit number you decrease its value by 63. find the number if the sum of its digits is 11.
Answer by lwsshak3(11628)

when you reverse the digits in a certain two digit number you decrease its value by 63. find the number if the sum of its digits is 11.
**
let u=units digit
let t=tens digit
..
t+u=11
number=10t+u
reverse number=10u+t
..
10u+t=10t+u-63
t=11-u
10u+11-u=10(11-u)+u-63
10u+11-u=110-10u+u-63
18u=36
u=2
t=11-u=9
10t+u=90+2=92
ans:
number:92